OpenStudy (anonymous):

A spherical ball is immersed in water contained in a vertical cylinder. The rise in water level is measured in order to calculate the radius of the ball.Calculate the radius of the ball in the following cases: (a): cylinder of radius 10cm, water level rises 4cm. (b): cylinder of radius 100cm, water level rises 8cm

OpenStudy (anonymous):

Here are the formulas for the cylinder and spheres respectively. \[V_{cyl}=\pi*r^2*h\] \[v_{sph}=\frac{4}{3}\pi*r^3\] you know the the sphere, when immersed, increases the volume. By how much? Well the volume of the cylinder without the sphere is V=pi*r^2*h The Volume of the cylinder afterwards is V=pi*r^2*(h+delta h), that means that the sphere must be the volume of the difference! \[v_{sph}=v_{cyl+\delta*h}-v_{cyl}=\pi*r^2*(h+\delta*h)-\pi*r^2*h=\pi*r^2*\delta*h=\pi*(10cm)^2*(4cm)\] \[v_{sph}=\frac{4}{3}\pi*r^3=\pi*(10cm)^2*(4cm)\] Now, you can solve for the radius. (b) is a similar case.